pts) Given a two locus system with genes A and B, where A1 is dominant to A2 and B1 is codominant to B2, what is the expected frequency of progeny exhibiting the A1 and B1 phenotypes only given the cross: A1A2B1B2 x A1A1B1B2?
Answer #1

In the case of dominant one allele suppress the effect of another gene while in codominance both the genes express itself as in the case of M and N blood types.

A1A2B1B2 x A1A1B1B2

i) monohybrid cross of A1A2 x A1A1

A1 A2
A1 A1A1 A1A2
A1 A1A1 A1A2

A1 is dominant to A2 so all the phenotypes exhibit A1 phenotype.

Frequency of A1 phenotype = 4/4 or 100℅

ii) monohybrid cross of B1B2 and B1B2

B1 B2
B1 B1B1 B1B2
B2 B1B2 B2B2

In B1B2 x B1B2, the phenotype ratio is

B1B1: B1B2: B2B2 = 1: 2: 1

B1 is codominant to B2 so genotype B1B2 exhibit both the phenotype

Frequency of B1 phenotype is 1/4

A1A2B1B2 x A1A1B1B2

The frequency of progeny exhibiting both A1 and B1 phenotypes only = 4/4 * 1/4 = 4/16 = 1/4 = 0.25 or 25%